# Class-Discriminative Attention Maps for Vision Transformers

### v1

@@ Abstract
Interpretability methods are critical components for examining and exploring deep neural networks (DNN), as well as increasing our understanding of and trust in them. Vision transformers (ViT), which can be trained to state-of-the-art performance with a self-supervised learning (SSL) training method [1], provide built-in attention maps (AM). While AMs can provide high-quality semantic segmentation of input images, they do not account for any signal coming from a downstream classifier. We introduce class-discriminative attention maps (CDAM), a novel post-hoc explanation method that is highly sensitive to the target class. Our method essentially scales attention scores by how relevant the corresponding tokens are for the predictions of a classifier head. Alternative to classifier outputs, CDAM can also explain a user-defined concept by targeting similarity measures in the latent space of the ViT. This allows for explanations of arbitrary concepts, defined by the user through a few sample images. We investigate the operating characteristics of CDAM in comparison with relevance propagation (RP) [2] and token ablation maps (TAM), an alternative to pixel occlusion methods. CDAM is highly class-discriminative and semantically relevant, while providing implicit regularization of relevance scores.
@@
Figure 1: Extending attention maps (AM), the proposed class-discriminative attention maps (CDAM) quantify and visualize input features that are relevant for the target class in Vision Transformers (ViT). We trained a linear classifier on the ViT-S/8 trained with DINO as a backbone [1] and visualize random samples with CDAM for different target classes. For details, refer to Section 3.1. Orange and blue colors correspond to positive and negative values, respectively.
@@ 1 Introduction
The transformer architecture, which was originally developed for natural language processing [3], has been developed into vision transformers (ViT) that demonstrate comparable or even better performance than convolutional neural networks (CNN) on many computer vision tasks [4, 5]. Additionally, a powerful feature of transformers is that the attention mechanism provides a built-in explanation of the model’s inner workings, if attention is understood as a weighing of the relevance of individual tokens [6, 7, 8, 9]. While attention maps (AM) provides intuitive interpretability, it fails to be discriminative with respect to a target class. Recognizing the potential of AM for explaining a ViT’s inner workings, we propose class-discriminative attention maps (CDAM), a gradient-based extension that makes them sensitive to class-specific signals coming from a downstream classifier.
@@
Recently, it has been demonstrated that ViT models trained with DINO, a self-supervised learning (SSL) training method [1], produce AMs that provide high-quality semantic segmentations of input images. These AMs clearly distinguish objects of interest from the background with very little noise, with different attention heads often focusing on semantically distinct regions [1]. AMs are the visualization of the self-attention of the [CLS] token query in the final attention layer. However, one shortcoming of AMs is that they purely operate on the level of the last attention layer and therefore do not take into account any information coming from a downstream task, such as a classifier head.
@@
The proposed CDAMs rely on the calculation of gradients of the class score, i.e. the logit value of a target class, with respect to the token activations in the final transformer block (Fig. 2(a)). We show through explicit calculation and empirical observations that CDAM scales the attention by how relevant the corresponding tokens are for the model’s decision. CDAM therefore inherits the desirable properties of AMs, but additionally indicates the (counter-)evidence for a target class that the ViT relies on for its predictions (Fig. 1). We demonstrate that it is class-discriminative in the sense that the heat maps clearly distinguish the object representing the target class from the background and, importantly, from objects that belong to other classes. CDAMs are also highly sensitive to the target class which, all together, makes it a useful tool for explaining the decision-making process of classifiers with a ViT backbone.
@@
In addition to the classifier-based explanations, we present a concept-based approach that replaces the class score with a similarity measure in the latent space of the ViT, where a concept is defined through a number of example images. Averaging the latent representations of the example images yields a concept embedding, which is analogous to a concept vector in variational autoencoders (VAE) and related generative models [10, 11]. In brief, we use CDAM to identify image patches that drive the (dis-)similarity of the concept embedding and the latent representation of a given image. This allows obtaining a heat map for the target concept the model has never been explicitly trained on, which is of particular interest for self-supervised models. Both supervised and semi-supervised variants of CDAM are compared to token ablation maps (TAMs) and relevance propagation (RP) [2]. Overall, our results suggest that CDAMs exhibit semantic coherency, implicit regularization, and strong discrimination with respect to a target class/concept.
@@ 2 Related Works
Interpretability of ViT, and more broadly deep neural networks (DNN), is an active area of research, since understanding their decision-making process would help improve the model, identify weaknesses and biases, and ultimately increase human trust in them. The proposed class-discriminative attention map (CDAM) is related to both attention and gradient-based explanation methods for deep learning models. On the one hand, explanations are obtained via backpropagation of gradients, but on the other hand, we show that the CDAM scores are directly proportional to the attention.
@@
Gradient-based methods operate by backpropagating gradients from the prediction score to the input features to produce feature attribution maps, also called saliency maps. Many variants have been proposed, see [12, 13, 14, 15, 16] to just name a few. In particular, CDAM is similar to and inspired by the input×$\times$×gradient method [16], because it involves calculating the gradients with respect to tokens multiplied by their activation (Eqs. 2 and 4). The main difference to the input×$\times$×gradient method, and most of other gradient methods, is that we don’t backpropagate the gradients all the way to the input tokens. CDAM stops backpropagation at the tokens that enter the final attention layer. In that regard, our method is related to GradCam [14] which backpropagates the gradients to the final convolutional feature map in a CNN. GradCam and CDAM therefore share the property that they operate on high-level features which is presumably the reason why they are both sensitive to the targeted class, a feature that many explanation methods do not have [17]. Several methods have attempted to apply gradient-based methods or Layer-Wise Relevance Propagation (LRP) [18] to Transformer architectures [19, 20, 21, 22, 23], mostly in the context of NLP tasks.
@@
Attention has often been used to gain insight into a model’s inner workings [6, 7, 8, 9] since it seems intuitive that this weighing of individual tokens correlates with how important they are for the performed task. There is also an ongoing discussion about whether attention provides meaningful explanations [8, 24, 25]. Furthermore, methods such as “attention rollout” [22] have been proposed that aim to quantify how the attention flows through the model to the input tokens. [2] defines the relevancy matrices, as the Hadamard product of the attention map and the gradient of $f_{c}$ w.r.t. to the attention layer, accumulated at each layer. This relevance propagation (RP) method [2], to the best of our knowledge, shows the highest degree of class-discrimination in ViT models. This is used in our study as a comparison for our method.
Many papers have been published on the problem of evaluating post-hoc explanation methods, leading even to the proposal of meta-evaluations [26]. In a recent overview of quantitative evaluation methods [27] 12 desirable properties have been identified. We primarily focus on three of them, namely correctness, contrastivity, and compactness. Our evaluation of correctness is a variant of Sensitivity-$n$ [28], which repeatedly masks $n$ randomly selected pixels and compares the sum of their relevance scores with the change in model output. In the context of ViTs, instead of arbitrarily perturbing observed pixels to some non-informative values, our method called token ablation map (TAM) removes patch tokens before they enter the first transformer block, followed by measuring the resulting difference in model performance. We consider TAM as an intuitive first-order approximation that serves as a sanity check for CDAM. Compactness is measured by looking at the shrinkage and sparsity of CDAM compared to other methods.
3 Proposed Methods
Figure 2: Graphical scheme of class-discriminative attention maps (CDAM). (a) CDAMs are obtained by first propagating an image through the transformer blocks and using its latent representation $l=\texttt{[CLS]}^{\prime}$ to infer a class score $f_{c}$ or similarity score $g(l,l_{c})$. The gradients wrt the layer-normalized input tokens $T_{i}$ of the final attention layer and their activations are then used to calculate the relevance score $S_{i,c}$ according to Eq. 2 or Eq. 4. (b) Detailed view of the final transformer block, adapted from [4].
@@ 3 Proposed Methods
In the first scenario, the class score is obtained from the corresponding activation in the prediction vector of a classifier trained on top of the ViT in a supervised fashion. In the second scenario for the concept similarity score, we define a concept through a small set of examples (e.g. ten images of a dog) and obtain a concept vector by averaging their latent representations. Then, without training a classifier, a similarity measure between the concept vector and the latent representation of a targeted sample image is computed and used as the target for the gradients. We call the resulting feature attribution maps “class-discriminative attention maps (CDAM)” because we observe that they scale the attention scores
@@
by how relevant the corresponding tokens $T_{i}$ are for the class or similarity score. $Q$ and $K$ are the query and key matrices, consisting of a set of $d_{k}$ dimensional vectors [3]. We discuss the proportionality of CDAM to these attention scores in Section 5.1.
3.1 Relevance Score from Classifiers
@@
where $T_{ij}$ is the $j$-th component of the $i$-th token that is fed to the last attention layer [^1] We found in our experiments that calculating the gradients wrt the layer-normalized tokens yields far better results than using the unnormalized ones. (Fig. 2). This definition is a first-order estimation of the contribution of $T_{i}$ to $f_{c}$ in the sense that it would be exact if $f_{c}$ depended on $T_{ij}$ only linearly. It is inspired by gradient-based explanation methods with the difference that we don’t backpropagate the gradients all the way to the input layer but only to the input tokens to the final attention layer. This allows us to gain insight into the decision-making process of the combination of ViT and classifier based on high-level features.
This approach leads to attention maps that are class-discriminative and well-aligned with the semantic relationships of the class and different parts of an image (Fig. 1). As a side note, the heatmaps shown in this article are all scaled independently, thus there is no meaningful comparison of the scale of relevance between images.
3.2 Relevance Score from Similarity Measures
@@
This approach is directly inspired by the concept saliency map [10]. The definition of the relevance score in this setting is analogous to Eq. 2
@@
Possible choices for $g$ are for example the $L^{2}$ distance, cosine similarity or dot product, where we use the latter in our experiments. The definition Eq. 4 estimates the contribution of $T_{i}$ to $g$, i.e. it allows us to gauge how much each token is driving the (dis-)similarity between $l$ and $l_{c}$.
3.3 Token Ablation Study
As an independent comparison method, we carried out a token ablation study, in which a change in the model accuracy is measured after removing one token at a time. We visualize changes in the model accuracy as token ablation maps (TAM). This method for ViT is analogous to input pixel occlusion [30]. Such occlusion methods gauge a (group of) pixel’s relevance by masking them with non-informative values and recording the change in model output. This approach is problematic because perturbing pixels potentially leads to out-of-distribution samples which may trigger artifacts of the model [31, 32]. Relatedly, in contrast to pixel occlusion methods, for TAM there is no need to choose a specific way of perturbing input features in pixel space, which would lead to some ambiguity and variation in the resulting relevance maps. The token ablation study sidesteps this issue in the context of ViTs by removing image patch tokens, one at a time, before they enter the first transformer block, instead of masking image patches in pixel space. The relevance score assigned by TAM is defined by
$S^{\text{TAM}}_{i,c}=f_{c}(\mathbf{T})-f_{c}(\mathbf{T}\backslash{T_{i}}),$ (5)
where $\mathbf{T}$ is the set of all tokens that enter the first transformer block and $\mathbf{T}\backslash{T_{i}}$ has token $T_{i}$ removed from $\mathbf{T}$.
Note that TAMs are not to be confused with the “ground truth” of feature relevance since we are dealing with non-linear models with complex interactions of patch tokens. Therefore, tokens would likely be assigned different relevance scores if we perturbed two or more tokens simultaneously, with different scores depending on the chosen group of tokens. This effect has previously been demonstrated in the context of CNNs [28]. Therefore, comparing with TAM mainly serves as a sanity check for CDAM, since TAM is a first approximation to patch token relevance with a clearly defined meaning.

### v2

@@ Abstract
Importance estimators are explainability methods that quantify feature importance for deep neural networks (DNN). In vision transformers (ViT), the self-attention mechanism naturally leads to attention maps, which are sometimes used as importance scores for which input features ViT models are focusing on. However, attention maps do not account for signals from downstream tasks. To generate explanations that are sensitive to downstream tasks, we have developed class-discriminative attention maps (CDAM), a gradient-based extension that estimates feature importance with respect to a known class or a latent concept. CDAM scales attention scores by how relevant the corresponding tokens are for the predictions of a classifier head. In addition to targeting the supervised classifier, CDAM can explain an arbitrary concept shared by selected samples by measuring similarity in the latent space of ViT. Additionally, we introduce Smooth CDAM and Integrated CDAM, which average a series of CDAMs with slightly altered tokens. Our quantitative benchmarks include correctness, compactness, and class sensitivity, in comparison to six other importance estimators. Vanilla, Smooth, and Integrated CDAM excel across all three benchmarks. In particular, our results suggest that existing importance estimators may not provide sufficient class-sensitivity. We demonstrate the utility of CDAM in medical images by training and explaining malignancy and biomarker prediction models based on lung Computed Tomography (CT) scans. Overall, CDAM is shown to be highly class-discriminative and semantically relevant, while providing compact explanations.
@@ 1 Introduction
Vision transformers (ViT) are a family of computer vision models based on the transformer architecture [1], that demonstrate comparable or even better performance than convolutional neural networks (CNN) on many object detection and classification tasks [2, 3]. A powerful feature of the transformers is that the attention mechanism provides a built-in explanation of the model’s inner workings, if attention is understood as a weighing of the relevance of individual tokens [4, 5, 6, 7]. While attention maps provide intuitive and inherent interpretability, they fail to be discriminative with respect to a target class. Recognizing the potential of attention maps for explaining a ViT’s inner workings, we have developed class-discriminative attention maps (CDAM), a gradient-based extension that makes them sensitive to class-specific signals based on a classifier or a concept derived from selected samples.
@@
Attention maps are the visualization of the self-attention of the class token query [CLS] in the final attention layer. Attention heads can be visualized separately, and in certain instances, may be related to semantically meaningful concepts (e.g., in language [8]). Recently, ViT trained with self-supervised learning (SSL) methods such as DINO [9] have shown excellent performance leveraging a greater amount of unlabeled data. Attention maps from such self-supervised ViT models provide high-quality semantic segmentations of input images. These attention maps clearly distinguish regions of interest (ROI) from the background with very little noise, with different attention heads often focusing on semantically distinct regions [9]. There have been many improvements to ViT including the introduction of convolution [10, 11] and spatial sparsity [12]. Attention maps, as well as our proposed CDAMs, are applicable to all ViT architectures with self-attention.
@@
The major shortcoming of attention maps, when considering them as an explainable AI (XAI) method, is that they purely operate on the level of the last attention layer and therefore do not take into account any information coming from a downstream task, such as a classifier head. The proposed CDAM overcomes this shortcoming by calculating the gradients of the class score, i.e. the logit value of a target class, with respect to the token activations in the final transformer block (visualized in Fig. 2(a)). Essentially, CDAM scales the attention scores by how relevant the corresponding tokens are for the model’s decision. CDAM therefore inherits the desirable properties of attention maps, but additionally indicates the evidence or counter-evidence for a target class that the ViT relies on for its predictions (examples in Fig. 1). We demonstrate that it is class-discriminative in the sense that the heat maps clearly distinguish the object representing the target class from the background and, importantly, from objects that belong to other classes. Therefore, CDAM is a useful tool for explaining the decision-making process of classifiers with a ViT backbone.
@@
In addition to the classifier-based explanations, we present a concept-based approach that replaces the class score with the similarity measure in the latent space of the ViT, where a concept is defined through selected images. Averaging the latent representations of the example images yields a concept embedding, which is analogous to a concept vector in variational autoencoders (VAE) and related generative models [13, 14]. This allows obtaining a heat map for the target concept the model has never been explicitly trained on, which is of particular interest for self-supervised models. We further demonstrate when computing gradients of a class (Section 3.1) or concept similarity (Section 3.2) score, with respect to the token activations in the final transformer block, smoothing or averaging could be applied to improve performance. Specifically, we introduce two additional variants of CDAM; (1) Smooth CDAM averages over multiple gradients that are computed with additional noise to tokens (inspired by SmoothGrad [15]) and (2) Integrated CDAM takes the intergral of the gradients along the path from the baseline to the token in the final transformer block (inspired by IntGrad [16]).
We conduct several quantitative evaluations focusing on correctness, class sensitivity, and compactness. Correctness is measured by perturbing an increasing amount of input features and measuring the model output [17, 18]. By using the ImageNet samples [19] with multiple objects [20] and applying importance estimators for different classes, we quantify the level of class-discrimination. Sparsity and shrinkage of explanations are measured by the number of near-zero importance scores. Besides vanilla, Smooth, and Integrated CDAM, we also employed attention maps, relevance propagation (RP) [21], partial Layer-wise Relevance Propagation (LRP) [22], Input $\times$ Gradients, SmoothGrad [15], and Integrated Gradient [16] in comparison. Three variants of CDAM outperform other explainability methods on three benchmarks.
Lastly, we have applied CDAM on a ViT fine-tuned on the Lung Image Database Consortium image collection (LIDC) [23]. Two ViT-based models are built for predicting malignancy and biomarkers. CDAM, as well as attention maps, are obtained. CDAM is shown to be highly class-discriminative and semantically relevant, while providing implicit regularization of importance scores.
@@ 2 Related Works
Interpretability of ViT, and more broadly explainable artificial intelligence (XAI), is an active area of research, since understanding their decision-making process would help improve the model, identify weaknesses and biases, and ultimately increase human trust in them. The proposed class-discriminative attention map (CDAM) is related to both attention and gradient-based explanation methods for deep learning models.
@@
Gradient-based methods operate by backpropagating gradients from the prediction score to the input features to produce feature attribution maps, which is also called saliency maps. Many variants have been proposed, see [24, 16, 25, 15, 26] to just name a few. In particular, CDAM is similar to and inspired by the input×$\times$×gradient method [26], because it involves calculating the gradients with respect to tokens multiplied by their activation (Eqs. 2 and 4). The main difference to most gradient-based methods is that we don’t backpropagate the gradients all the way to the input tokens. CDAM stops backpropagation at the tokens that enter the final attention layer. In that regard, our method is related to GradCam [25] which backpropagates the gradients to the final convolutional feature map in a CNN. GradCam and CDAM therefore share the property that they operate on high-level features which is presumably the reason why they are both sensitive to the targeted class, a feature that many explanation methods do not have [27]. Several methods have attempted to apply gradient-based methods or Layer-Wise Relevance Propagation (LRP) [28] to Transformer architectures [29, 30, 22, 31, 32], mostly in the context of NLP tasks.
@@
Attention has often been used to gain insight into a model’s inner workings [4, 5, 6, 7] since it seems intuitive that this weighing of individual tokens correlates with how important they are for the performed task. There is also an ongoing discussion about whether attention provides meaningful explanations [6, 33, 34]. Furthermore, methods such as “attention rollout” [31] have been proposed that aim to quantify how the attention flows through the model to the input tokens. [21] defines the relevancy matrices, as the Hadamard product of the attention map and the gradient of $f_{c}$ w.r.t. to the attention layer, accumulated at each layer. This relevance propagation (RP) method [21], to the best of our knowledge, shows the highest degree of class-discrimination in ViT models. This is used in our study as one of comparative importance estimators.
Many papers have been published on the problem of evaluating post-hoc explanation methods, leading even to the proposal of meta-evaluations [35]. In a recent overview of quantitative evaluation methods [36] 12 desirable properties have been identified. We primarily focus on three of them, namely correctness, contrastivity, and compactness. In particular, carefully designed occulusion/perturbation studies can help us calculate a fidelity statistics [17], which is also called the symmetric relevance gain (SRG) measure [37]. [38] first demonstrated that many importance estimators, which purport to be sensitive to the internals of a classifier, produce similar importance scores regardless of the relation of inputs and classes in the training data, which they called “the data randomization test”. Generally, if two importance scores with respect to two different classes produce very similar heat maps, the importance estimator is not class sensitive. We quantify this concept in our class sensitivity evaluation benchmark to measure whether an explanation is sensitive to the classes.
3 Methods
Adaptation of the transformer architecture to computer vision hinges on creating tokens from small patches of the input image [2]. In addition to those image patch tokens, the [CLS] token is added, which is utilized for downstream tasks such as classification. Attention maps are the two-dimensional visualization of the self-attention in the final attention layer with [CLS] as query token, where the attention scores are averaged over all heads.
@@ 3 Methods
In the first scenario, the class score is obtained from the corresponding activation in the prediction vector of a classifier trained on top of the ViT in a supervised fashion. In the second scenario, we define a concept through a small set of examples (e.g. ten images of a dog) and obtain a concept vector by averaging their latent representations. Then, without training a classifier, a similarity measure between the concept vector and the latent representation of a targeted sample image is computed and used as the target for the gradients. We call the resulting maps of feature relevance “class-discriminative attention maps” because they scale the attention scores
@@
by how relevant the corresponding tokens $T_{i}$ are for the class or similarity score. $Q$ and $K$ are the query and key matrices, consisting of a set of $d_{k}$ dimensional vectors [1]. We discuss the proportionality of CDAM to the attention scores in Section A.1.
Self-supervised learning (SSL) methods such as DINO can train ViT models whose attention maps are high-quality semantic segmentations of the input, without using labels [9]. Therefore, we utilize a ViT model [Dosovitskiy2020ViT] pre-trained with DINO as a backbone for our experiments and applications. We denote details about the ViT architecture, training schemes, and model performance in the Section 4. Note that the heat maps shown in this article are scaled independently.
3.1 CDAM for a known class
@@
where $T_{ij}$ is the $j$-th component of the $i$-th token that is fed to the last attention layer (Fig. 2). We found in our experiments that calculating the gradients with respect to the layer-normalized tokens yields far better results than using the unnormalized ones.
This definition is a first-order estimation of the contribution of $T_{i}$ to $f_{c}$ in the sense that it would be exact if $f_{c}$ depended on $T_{ij}$ only linearly. It is inspired by gradient-based explanation methods with the difference that we don’t backpropagate the gradients all the way to the input layer but only to the input tokens to the final attention layer. This allows us to gain insight into the decision-making process of the combination of ViT and classifier based on high-level features. This approach leads to attention maps that are class-discriminative and well-aligned with the semantic relationships of the class and different parts of an image (Fig. 1).
3.2 CDAM for a latent concept
@@
This approach is directly inspired by the concept saliency map [13]. The definition of the importance score in this setting is analogous to Eq. 2
@@
Possible choices for $g$ are for example the $L^{2}$ distance, cosine similarity or dot product, where we use the latter in this article. The definition Eq. 4 estimates the contribution of $T_{i}$ to $g$, such that it allows us to gauge how much each token is driving the similarity between $l$ and $l_{c}$.
3.3 Smooth and Integrated CDAM
SmoothGrad (SG) [15] and Integrated Gradients (IG) [16] extended saliency maps [24], which were introduced in the context of CNNs as visualizing gradients of the class with respect to the input features. Essentially, these two modifications sharpen and denoise vanilla saliency maps, with theoretical and practical justifications. Briefly, [15] adds a small Gaussian noise to a sample image $n$ times, resulting in $n$ vanilla saliency maps. SG is an average of those $n$ saliency maps. In IG, instead of adding random noise, [16] uses multiple saliency maps obtained from interpolating between a given sample and a baseline (e.g., a black image).
We have adapted the analogous modifications to CDAM, resulting in two additional variants which we call Smooth and Integrated CDAM. Instead of modifying the input images, we vary the tokens that enter the final transformer block. In Integrated CDAM, we interpolate between a baseline and the tokens themselves, resulting in $n$ maps. In Smooth CDAM, we repeatedly add a small noise to the tokens and again obtain $n$ maps.
Specifically, Smooth and Integrated CDAM are defined as follows:
$S_{i,c}^{\text{SG}}=\frac{1}{n}\sum_{1}^{n}\sum_{j}T^{\prime}_{ij}\frac{\partial f_{c}(T^{\prime})}{\partial T^{\prime}_{ij}}$ (5)
$S_{i,c}^{\text{IG}}=\sum_{j}(T_{ij}-\tilde{T}_{ij})\times\frac{1}{n}\sum_{1}^{n}\frac{\partial f_{c}(\tilde{T}+k/n\times(T-\tilde{T}))}{\partial T_{ij}}$ (6)
where $T_{i}=T_{i}+\mathcal{N}\left(0,\sigma^{2}\right)$ with $\mathcal{N}\left(0,\sigma^{2}\right)$ representing Gaussian noise with standard deviation $\sigma$. For the baseline $\tilde{T}_{i}$ we choose the null vector. We implement these two variants by first propagating an image through the ViT up to the final transformer block, where the activations of the tokens are obtained. Next, the activations are manipulated by either applying the noise or interpolation step and the manipulated tokens are then further propagated to obtain the final ViT output.
3.4 Quantitative Evaluation
We perform quantitative evaluations of the correctness, class sensitivity, and compactness of importance estimators. In addition to our proposed methods (vanilla, smooth, and integrated CDAM), we perform evaluations on the raw attention maps, relevance propagation (RP) [21], partial Layer-wise Relevance Propagation (LRP) [22], SmoothGrad [15], Integrated Gradient [16], and Input $\times$ Gradients. Note that based on classic LRP [28], partial LRP [22] provides improvement for multi-head self-attention mechanisms in modern transformer architecture [22]. Relevance propagation is a state-of-the-art importance estimator for explaining a prediction (e.g., class discriminatory) based on ViT [21]. We use 50 interpolation steps for Integrated Gradient and Integrated CDAM and chose the standard deviation of the noise used in SmoothGrad and smooth CDAM based on the best performance on the fidelity benchmark (explained below) on a held-out validation set of 200 images.
To evaluate the correctness (also known as faithfulness) of importance scores, we employ two versions of input perturbation-based approaches [40, 41, 42]. Generally, an input feature (e.g., a pixel) is masked and a resulting model performance is observed. In all cases, input features are perturbed by being replaced with their blurred versions. A Gaussian blur with $\sigma=14$ turns pixels uninformative while minimizing perturbation artifacts [17].
Fidelity. The first evaluation method uses accuracy curves obtained from the Most Important First (MIF) and the Least Important First (LIF) perturbations, which are abbreviated as MIF/LIF perturbation [17, 18]. Pixels are ranked by importance scores from a specific explainability method (e.g., CDAM) and are perturbed in the MIF or LIF order. In the MIF perturbation, an accurate importance score should lead to a relatively rapid decrease of the model output (logit) for the target class. For the LIF perturbation, we expect an inverse relationship. Note that some importance scores are unsigned (i.e., only non-negative values) and others are signed. For signed importance, with negative scores faithfully indicating counter-evidence, there may be an increase in the logits if pixels with negative scores are masked. For unsigned and correct importance scores, one expects initially constant logits in the LIF perturbation.
The model output $f_{\text{MIF}}(p)$ or $f_{\text{LIF}}(p)$ as a function of the fraction of perturbed pixels $p$ gives the perturbation curve [17, 18]. We average these curves over all samples, providing an overview of fidelity for a given importance estimator. We further summarize the MIF and LIF perturbation evaluation by using the area $A_{\text{LIF-MIF}}$ (called the fidelity statistics) under the difference of perturbation curves $f_{\text{LIF}}-f_{\text{MIF}}$ to quantify the fidelity of the explanation methods [18]. Generally speaking, a higher value of fidelity statistics $A_{\text{LIF-MIF}}$ indicates a better explanation [17].
Box sensitivity. For the second evaluation method, inspired by Sensitivity-n [43], we measure how the model output (logit) changes as a group of pixels are perturbed. Pixels are selected by a randomly placed square box (of size $s\times s$); then, the sum of their importance scores and the change in the model output are measured. For accurate importance estimators, we expect to see a high correlation between the two measures.
For a given image, we sample these values 100 times and calculate the Pearson correlation between the sum of importance scores and the model output change. This is performed for a range of differently sized boxes to obtain $f_{\text{box}}(s)$, the correlation statistics as a function of the box size $s$. These sensitivity curves are averaged over all samples. Performance of importance estimators varies with $s$, and therefore we use the area $A_{\text{box}}$ under the box sensitivity curve to quantify correctness. Generally, a larger box sensitivity statistics $A_{\text{box}}$ implies greater correctness.
In contrast to Sensitivity-n which randomly selects and perturbs $n$ pixels [43], we use a random square of pixels. A box removes a certain region which mimics how image patches are processed by ViT. Box sensitivity may therefore remove larger parts or complete objects. In contrast, removing random pixels may leave enough neighboring pixels with sufficient semantic information such that perturbation may not have any impact of the model output. Indeed, we have observed that for Sensitivity-n, the correlation is very low ($<0.1$) for all methods.
Class discrimination. We have designed an evaluation benchmark to measure class discrimination, which can be applied on images containing multiple objects (see $X_{\text{multi}}$ dataset below). We obtain explanations for two distinct classes corresponding to the present objects. When performing the aforementioned evaluation benchmarks (fidelity and box sensitivity), we measure the change in the model output with respect to one of the classes. Evaluation metrics are thus obtained using importance scores based on either the correct target class (denoted by MIF/LIF or box) or a wrong class (denoted by MIF/LIF or box).
For important estimators that are not sensitive/discriminative to the target class, the importance scores are similar (or the same) regardless of which class is targeted. The change of the model output during perturbation would remain the same, irrespective of the class that was targeted to obtain the importance ranking of pixels. For highly class-discriminative importance estimators, importance scores would be different depending on which class is targeted. Furthermore, for accurate class-discriminative methods, the performances on fidelity and box sensitivity benchmarks can be expected to deteriorate substantially when perturbing according to the importance scores targeting the wrong class.
The areas under MIF/LIF perturbation or box sensitivity curves, when using a wrong class, are denoted by $A_{\text{LIF}^{\prime}-\text{MIF}^{\prime}}$ and $A_{\text{box}^{\prime}}$, respectively. The areas $\Delta A_{\text{LIF-MIF}}$ and $\Delta A_{\text{box}}$ under the respective perturbation curves
$\Delta f_{\text{LIF-MIF}}$ $=(f_{\text{LIF}}-f_{\text{MIF}})-(f_{\text{LIF}^{\prime}}-f_{\text{MIF}^{\prime}}),$ (7)
$\Delta f_{\text{box}}$ $=f_{\text{box}}-f_{\text{box}^{\prime}}.$ (8)
measure the degree of class-discrimination. In particular, if an importance estimator is not sensitive to the target class at all, $\Delta f_{\text{LIF-MIF}}\approx\Delta f_{\text{box}}\approx 0$. An importance estimator that performs poorly in the data randomization test of [38] would have low $\Delta A_{\text{LIF-MIF}}$ and $\Delta A_{\text{box}}$.
Compactness. We evaluate sparsity and shrinkage by counting the number of input pixels with importance scores smaller than $5\%$ of the maximum value. A threshold is defined by $t\times S_{c}^{\text{max}}$ with $t=0.05$, where $S_{c}^{\text{max}}$ is the maximum importance score for a given image and an importance estimator. Generally, a compact explanation is considered desirable, where sparsity and shrinkage are quantitative approaches [44].
3.5 Data
We perform quantitative evaluations on two datasets. The first one, $X_{\text{val}}$, is a randomly selected subset of the ImageNet validation set consisting of 1000 images [19]. The second one, $X_{\text{multi}}$, selects 1000 images from the validation set that contain multiple objects [20]. We filtered out instances in which the class labels provided by [20] actually refer to the same object (e.g. notebook and laptop) by calculating and thresholding the cosine similarity of the embedded class names. Embeddings were obtained using OpenAI’s text-embedding-3-large model.
The $X_{\text{multi}}$ dataset is used to evaluate the degree of class-discrimination exhibited by the explanations. To this end, we obtain explanations w.r.t to two annotated classes, perform the MIF/LIF perturbation and box sensitivity, and record the model output for the chosen target class. For example, for the top right sample in Fig. 1, we would obtain the explanations for the classes Tench and Golden Retriever (not shown) and record the model output for the class Golden Retriever. For the right target (Tench), we obtain a fidelity statistics $A_{\text{LIF}-\text{MIF}}$. For the wrong target (Golden Retriever), $A_{\text{LIF}^{\prime}-\text{MIF}^{\prime}}$. In the worst case scenario, where explanations for those two distinct classes are the same, $A_{\text{LIF}^{\prime}-\text{MIF}^{\prime}}=A_{\text{LIF}-\text{MIF}}$ and the resulting $\Delta A_{\text{LIF-MIF}}$ is zero. Equivalent for $\Delta A_{\text{box}}$.
For application in medical images, we use lung computed tomography (CT) scans that were collected from and validated by seven academic centers and eight medical imaging companies [23]. The Lung Image Database Consortium image collection (LIDC) [23] contains 1018 records of CT scans. Presence of nodules (nodule $\leq$ 3 mm; nodule $<$ 3 mm; non-nodule $\geq$ 3 mm) is annotated by up to 4 human annotators. Segmentation of nodules (i.e., ROI) is the average contour given by annotators. Eight biomarkers that are informative in clinical practices are also collected in LIDC: subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter. For clinical descriptions of biomarkers, refer to [45]. After preprocessing, 854 nodule crops of size, corresponding biomarkers, and labels (benign or malignant) were saved and used for downstream procedures. Overall, the LIDC dataset used in this study contains 443 benign examples and 411 malignant examples.